### v1

@@ 1 Introduction
$\mathrm{EMD}$, also known as geometric transportation or geometric bipartite matching, is a widely studied distance measure (see, e.g., [30, 7, 6, 8, 40, 32, 26, 25, 1, 3]) that has received significant interest in computer vision, starting with the work of [42]. Depending on the precise formulation, $\mathrm{EMD}$ is a distance measure on point sets, distributions, or functions. In this paper, we study the following formulation of $\mathrm{EMD}$ as measuring the distance from a set of blue points $B$ to a set of red points $R$:
@@
We briefly discuss algorithms for $\mathrm{EMD}$. Note that $\mathrm{EMD}$ can be formulated as a mincost matching problem on a bipartite graph with vertices $R\cup B$, where edge lengths are equal to the point-to-point distances. This graph has $|R|\cdot|B|=\mathcal{O}(n^{2})$ edges and solving bipartite mincost matching by the Hungarian method yields an exact algorithm for $\mathrm{EMD}$ with running time $\mathcal{O}(n^{3})$. Alternatively, by combining geometric spanners with recent advancements in (approximate) mincost flow solvers, one can obtain fast approximation algorithms for $\mathrm{EMD}$. For instance, symmetric $\mathrm{EMD}$ in $L_{2}$ metric can be solved in time $n(\log(n)/\varepsilon)^{\mathcal{O}(d)}$ [32]. See also [30, 8, 26, 25, 1, 3] for more approximation algorithms. Conditional lower bounds are also known, but they apply only when the dimension is super-constant [40].
@@
Here, $B+\tau=\{b+\tau\mid b\in B\}$ is the translated point set. See Figure 1.1 for an illustration of this distance measure. Again, we call asymmetric $\mathrm{EMDuT}_{p}$ the problem of computing $\mathrm{EMDuT}_{p}(B,R)$ for given sets $B,R$ of size $|B|\leq|R|=n$, and the symmetric variant comes with the additional restriction $|B|=|R|$. This measure was introduced by Cohen and Guibas [19], who presented heuristics as well as an exact algorithm with respect to the squared Euclidean distance. Later, Klein and Veltkamp [33] designed a 2-approximation algorithm for symmetric $\mathrm{EMDuT}_{p}$ running in asymptotically the same time as any $\mathrm{EMD}$ algorithm. Cabello, Giannopoulos, Knauer, and Rote [14] designed $(1+\varepsilon)$-approximation algorithms for $\mathrm{EMDuT}_{2}$ in the plane, running in time $\widetilde{\mathcal{O}}(n^{4}/\varepsilon^{4})$ for the asymmetric variant and $\widetilde{\mathcal{O}}(n^{3/2}/\varepsilon^{7/2})$ for the symmetric variant. [^2] Here and throughout the paper we use $\widetilde{\mathcal{O}}$ notation to ignore logarithmic factors, i.e., $\widetilde{\mathcal{O}}(T)=\bigcup_{c\geq 0}\mathcal{O}(T(\log T)^{c})$. Eppstein et al. [23] proposed algorithms to solve the symmetric $\mathrm{EMDuT}_{1}$ and symmetric $\mathrm{EMDuT}_{\infty}$ problems in the plane, that run in $\mathcal{O}(n^{6}\log^{3}n)$ time. We remark that most of these works also study variants of $\mathrm{EMDuT}$ under more general transformations than translations, but in this paper we focus on translations.
We are not aware of any other research on $\mathrm{EMDuT}$, which is surprising, since translation-invariant distance measures are well motivated, and the analogous Hausdorff distance under translation [27, 41, 28, 2, 35, 34, 13, 16] and Fréchet distance under translation [4, 37, 31, 9, 11, 24, 12] have received considerably more attention.
@@ 1.1 Our results
We extend the work of Eppstein et al. [23] for point sets in $\mathbb{R}^{d}$, leading to the following algorithms.
@@
Given sets $B,R\subseteq\mathbb{R}^{d}$ of size $m=|B|\leq n=|R|$, $\mathrm{EMDuT}_{1}(B,R)$ and $\mathrm{EMDuT}_{\infty}(B,R)$ can be computed in time $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$.
We explain that such a dependence on dimension is unavoidable, by establishing a more coarse-grained lower bound compared to our 1D results: We show that no algorithm can solve the problem in time $n^{o(d)}$. In fact, we present a stronger lower bound that even rules out fast approximation algorithms. Our lower bound assumes the Exponential Time Hypothesis (ETH) [29], which is a well-established conjecture from fine-grained complexity theory.
@@ 1.2 Open problems
The $L_{2}$ metric is the most natural measure in the geometric settings, making $\mathrm{EMDuT}_{2}$ a well motivated problem. The most pressing open problem is to determine the complexity of the $\mathrm{EMDuT}_{2}$ problem in any dimension $d\geq 2$.
It is known that the $\mathrm{EMDuT}_{2}$ problem cannot be solved exactly. Namely, for any point set $R\subset\mathbb{R}^{d}$ of size $n$, if $B$ consists of $n$ copies of the point $(0,\ldots,0)$, then $\mathrm{EMDuT}_{2}(B,R)$ is the (cost of the) Geometric Median of $R$. Because the Geometric Median has no exact algebraic expression (even for $d=2$[10], there is no exact algorithm for $\mathrm{EMDuT}_{2}$ in dimension $d\geq 2$.
We therefore need to relax the goal and ask for an approximation algorithm. Geometric Median has a very fast $(1+\varepsilon)$-approximation algorithm running in time $\mathcal{O}(nd\log^{3}(1/\varepsilon))$ [18], so the reduction from Geometric Median to $\mathrm{EMDuT}_{2}$ does not rule out very fast approximation algorithms for $\mathrm{EMDuT}_{2}$.
@@ 3 Algorithm in one dimension
When $B,R\subset\mathbb{R}$ there is an optimal matching $\phi$ that is monotonically increasing.
@@
Assume we have an optimal matching $\phi$ where $\phi(b)=r^{\prime}$ and $\phi(b^{\prime})=r$ for $b>b^{\prime}$ and $r>r^{\prime}$. Then, define a modified matching $\tilde{\phi}$ that has $\tilde{\phi}(b)=r$ and $\tilde{\phi}(b^{\prime})=r^{\prime}$, and everywhere else is the same as $\phi$. Without loss of generality, assume $b\geq r$. Then, $|b-r^{\prime}|+|b^{\prime}-r|=|b-r|+|r-r^{\prime}|+|b^{\prime}-r|>|b-r|+|b^{\prime}-r^{\prime}|$ which means $\mathcal{D}_{B,R}(\phi,0)>\mathcal{D}_{B,R}(\tilde{\phi},0)$. Thus, we can modify the matching until it is monotonically increasing. ∎
@@ 3.1 Symmetric case
In the symmetric case ($|R|=|B|$), Lemma 3.1 uniquely defines an optimal matching. Let $B=\{b_{1},\ldots,b_{n}\}$ and $R=\{r_{1},\ldots,r_{n}\}$ be the points in increasing order. Now, the optimal translation $\tau^{*}$ is the value for $\tau$ that minimizes $\mathcal{D}_{B,R}(\phi,\tau)=\sum_{i=1}^{n}|b_{i}-r_{i}+\tau|$. Thus, it corresponds to the median of $b_{1}-r_{1},\ldots,b_{n}-r_{n}$, which we can compute in $O(n\log n)$ time.
@@ 3.2 Asymmetric case
In this section, we present an $\mathcal{O}(mn(\log n+\log^{2}m))$ time algorithm to compute $\mathrm{EMDuT}(B,R)$, for the case that $m\leq n$. Consider the cost $f(\tau)=\min_{\phi\in\Phi}\mathcal{D}_{B,R}(\phi,\tau)$ as a function of $\tau$. The minimum of this function is $\mathrm{EMDuT}(B,R)$. The main idea is then to sweep over the domain of $f$, increasing $\tau$ from $\infty$- ∞ to $\infty$, while maintaining (a representation of) $f$ and a matching $\phi$ that realizes cost $f(\tau)=\mathcal{D}_{B,R}(\phi,\tau)$. We also maintain the best translation $\tau^{*}\leq\tau$ (i.e. with minimal cost) among the translations considered so far (and if there are multiple such translations, the smallest one), so at the end of our sweep, $\tau^{*}$ is thus an optimal translation.
@@
By Lemma 3.1, for any $\tau$, there exists an optimal monotonically increasing matching between $B+\tau$ and $R$. So, we restrict our attention to such monotonically increasing matchings. Observe that any such matching $\phi$ corresponds to a partition of $B$ into runs, i.e. maximal subsequences of consecutive points, $B_{1},\ldots,B_{z}$, so that the points $b_{t-k},\ldots,b_{t}$ in a run $B_{i}$ are matched to consecutive red points $r_{u-k},\ldots,r_{r}$, for some $r_{u}=\phi(b_{t})$. Moreover, for any such a matching $\phi$, the function $\mathcal{D}_{B,R}(\phi,\tau)$ is piecewise linear in $\tau$, and each breakpoint is a translation $\tau$ for which there is a pair $(b,r)\in B\times R$ with $b+\tau=r$. It then follows that $f(\tau)$ is also piecewise linear in $\tau$. Furthermore, the breakpoints of $f$ are of two types. A type (i) breakpoint is a translation such that there is a pair $(b,r)\in B\times R$ with $b+\tau=r$, and a type (ii) breakpoint if there are two different matchings $\phi,\phi^{\prime}$ that both realize the same minimum cost $\mathcal{D}_{B,R}(\phi,\tau)=\mathcal{D}_{B,R}(\phi^{\prime},\tau)$. We show the following key lemma, which lets us characterize the breakpoints of type (ii) more precisely.
Figure 3.1: Schematic representation of the graph $G=\phi\otimes\phi^{\prime}$ used in the proof of Lemma 3.3. Each edge exists if and only if exactly one edge from either $\phi$ or $\phi^{\prime}$ is present. Green edges arise from the matching $\phi^{\prime}$, while yellow edges arise from the matching $\phi$.
@@
Let $\phi$ be an optimal monotone matching of $\mathrm{EMD}_{p}(B+\tau,R)$, and let ϕsuperscriptitalic-$\phi^{\prime}$ be an optimal monotone matching of $\mathrm{EMD}_{p}(B+\tau^{\prime},R)$ for some $\tau^{\prime}>\tau$. Then, $\phi^{\prime}(b)\geq\phi(b)$ for all $b\in B$.
@@
We assume that all the points are distinct, as otherwise we can perturb them infinitesimally to resolve ties. Consider a bipartite graph $G$ defined as follows: the vertices of $G$ are $B\uplus R$, and we add an edge $(b,r)\in B\times R$ to $G$ if exactly one of the following conditions holds: (i) $\phi(b)=r$ and $\phi^{\prime}(b)\neq r$, or (ii) $\phi^{\prime}(b)=r$ and $\phi(b)\neq r$. The graph $G$ can be thought of as the exclusive-or of the matchings $\phi$ and $\phi^{\prime}$, see Figure 3.1. We will now demonstrate that the connected components of this graph are paths. Then, considering that the matchings are monotone, it follows that the edges of these paths are non-crossing. This implies that consecutive red vertices on these paths are monotone. Hence, if we translate to the right, the matching $\phi^{\prime}$ is superior to $\phi$.
@@
Observe that the maximum degree of graph $G$ is $2$, so the connected components of $G$ consist of cycles and paths. Assume that $C$ is a cycle. In that case, however, there exists a pair of edges from either $\phi$ or $\phi^{\prime}$ that intersect (see Figure 3.2). This contradicts the assumption about the monotonicity of both $\phi$ and $\phi^{\prime}$. Hence, $C$ is not a cycle, and the proof of the claim follows. ∎
@@
Assume that $p_{1}$ is in $B$ (the proof for $p_{\ell}$ is analogous). If $\phi(b^{\prime})=\phi(b)$, then the vertex $p_{1}$ is isolated in $G$. If $\phi(b^{\prime})\neq\phi(b)$, then the degree of $p_{1}$ is $2$. Hence, $p_{1}$ cannot have a degree $1$. ∎
@@
In particular, there exists an index $k\in[\ell-2]$ such that $t_{k}\cdot t_{k+1}<0$. This means that either (a) $p_{k}>p_{k+2}$ and $p_{k+1}<p_{k+3}$, or (b) $p_{k}<p_{k+2}$ and $p_{k+1}>p_{k+3}$. In both of these cases, the intervals $(p_{k},p_{k+1})$ and $(p_{k+2},p_{k+3})$ intersect (see Figure 3.3), which contradicts the assumption about the monotonicity of $\phi$ and $\phi^{\prime}$. ∎
Figure 3.3: Two cases of Claim 3.6 in which the crossing occurs.
Now, we continue the proof of Lemma 3.3. For the sake of contradiction, assume that $\phi^{\prime}(b)<\phi(b)$ for some $b\in B$. By Claim 3.6, it actually means that $\phi^{\prime}(b)<\phi(b)$ for every $b\in B$. Recall that the cost of matching $\phi$ is:
$\mathcal{D}_{B,R}(\phi,\tau)=\sum_{b\in B}\lvert b+\tau-\phi(b)\rvert.$

### v2

@@ 1 Introduction
$\mathrm{EMD}$, also known as geometric transportation or geometric bipartite matching, is a widely studied distance measure (see, e.g., [Indyk07, AndoniIK08, AndoniBIW09, AndoniNOY14, Rohatgi19, KhesinNP20, FoxL22, FoxL23, AgarwalCRX22, AgarwalRSS22]) that has received significant interest in computer vision, starting with the work of [RubnerTG00]. Depending on the precise formulation, $\mathrm{EMD}$ is a distance measure on point sets, distributions, or functions. In this paper, we study the following formulation of $\mathrm{EMD}$ as measuring the distance from a set of blue points $B$ to a set of red points $R$:
@@
We briefly discuss algorithms for $\mathrm{EMD}$. Note that $\mathrm{EMD}$ can be formulated as a mincost matching problem on a bipartite graph with vertices $R\cup B$, where edge lengths are equal to the point-to-point distances. This graph has $|R|\cdot|B|=\mathcal{O}(n^{2})$ edges and solving bipartite mincost matching by the Hungarian method yields an exact algorithm for $\mathrm{EMD}$ with running time $\mathcal{O}(n^{3})$. Alternatively, by combining geometric spanners with recent advancements in (approximate) mincost flow solvers, one can obtain fast approximation algorithms for $\mathrm{EMD}$. For instance, symmetric $\mathrm{EMD}$ in $L_{2}$ metric can be solved in time $n(\log(n)/\varepsilon)^{\mathcal{O}(d)}$ [KhesinNP20]. See also [Indyk07, AndoniNOY14, FoxL22, FoxL23, AgarwalCRX22, AgarwalRSS22] for more approximation algorithms. Conditional lower bounds are also known, but only when the dimension is super-constant [Rohatgi19].
@@
Here, $B+\tau=\{b+\tau\mid b\in B\}$ is the translated point set. See Figure 1.1 for an illustration of this distance measure. Again, we call asymmetric $\mathrm{EMDuT}_{p}$ the problem of computing $\mathrm{EMDuT}_{p}(B,R)$ for given sets $B,R$ of size $|B|\leq|R|=n$, and the symmetric variant comes with the additional restriction $|B|=|R|$. This measure was introduced by Cohen and Guibas [CohenG99], who presented heuristics as well as an exact algorithm with respect to the squared Euclidean distance. Later, Klein and Veltkamp [KleinV05] designed a 2-approximation algorithm for symmetric $\mathrm{EMDuT}_{p}$ running in asymptotically the same time as any $\mathrm{EMD}$ algorithm. Cabello, Giannopoulos, Knauer, and Rote [CabelloGKR08] designed $(1+\varepsilon)$-approximation algorithms for $\mathrm{EMDuT}_{2}$ in the plane, running in time $\widetilde{\mathcal{O}}(n^{4}/\varepsilon^{4})$ for the asymmetric variant and $\widetilde{\mathcal{O}}(n^{3/2}/\varepsilon^{7/2})$ for the symmetric variant. [^1] Here and throughout the paper we use $\widetilde{\mathcal{O}}$ notation to ignore logarithmic factors, i.e., $\widetilde{\mathcal{O}}(T)=\bigcup_{c\geq 0}\mathcal{O}(T(\log T)^{c})$. Eppstein et al. [eppstein15improv_grid_map_layout_point_set_match] proposed algorithms to solve the symmetric $\mathrm{EMDuT}_{1}$ and symmetric $\mathrm{EMDuT}_{\infty}$ problems in the plane, that run in $\mathcal{O}(n^{6}\log^{3}n)$ time. We remark that most of these works also study variants of $\mathrm{EMDuT}$ under more general transformations than translations, but in this paper we focus on translations.
We are not aware of any other research on $\mathrm{EMDuT}$, which is surprising, since translation-invariant distance measures are well motivated, and the analogous Hausdorff distance under translation [HuttenlocherK90, Rote91, HuttenlocherRK92, AgarwalHSW10, KnauerS11, KnauerKS11, BringmannN21, Chan23] and Fréchet distance under translation [AltKW01, MosigC05, JiangXZ08, AvrahamKS15, BringmannKN20, FiltserK20, BringmannKN21] have received considerably more attention.
@@ 1.1 Our Results
We extend the work of Eppstein et al. [eppstein15improv_grid_map_layout_point_set_match] for point sets in $\mathbb{R}^{d}$, leading to the following algorithms.
@@
Given sets $B,R\subseteq\mathbb{R}^{d}$ of size $m=|B|\leq n=|R|$, $\mathrm{EMDuT}_{1}(B,R)$ and $\mathrm{EMDuT}_{\infty}(B,R)$ can be computed in $\mathcal{O}(m^{d}n^{d+2}\log^{d+2}n)$ time.
We explain that such a dependence on the dimension is unavoidable, by establishing a more coarse-grained lower bound compared to our 1D results: We show that no algorithm can solve the problem in time $n^{o(d)}$. In fact, we present a stronger lower bound that even rules out fast approximation algorithms. Our lower bound assumes the Exponential Time Hypothesis (ETH) [ImpagliazzoP01], which is a well-established conjecture from fine-grained complexity theory.
@@ 1.2 Open Problems
The $L_{2}$ metric is the most natural measure in geometric settings, making $\mathrm{EMDuT}_{2}$ a well motivated problem. The most pressing open problem is to determine the complexity of the $\mathrm{EMDuT}_{2}$ problem in any dimension $d\geq 2$.
We first observe that the $\mathrm{EMDuT}_{2}$ problem cannot be solved exactly (on the Real RAM model of computation). Indeed, on the Real RAM supporting only the usual arithmetic operations ($+,-,\cdot,/$), if the input numbers are rational then all output numbers are rational. If the machine further supports square roots (or other integral roots), if the input numbers are rational then the output numbers are algebraic. For the Geometric Median problem there are instances with rational input coordinates where the coordinates of the geometric median are not algebraic; this even holds in two dimensions [Bajaj88]. Therefore, Geometric Median cannot be solved exactly on the Real RAM. Finally, note that Geometric Median is a special case of $\mathrm{EMDuT}_{2}$, as for any point set $R\subset\mathbb{R}^{d}$ of size $n$, if $B$ consists of $n$ copies of the point $(0,\ldots,0)$, then $\mathrm{EMDuT}_{2}(B,R)$ is the (cost of the) Geometric Median of $R$. Therefore, also $\mathrm{EMDuT}_{2}$ cannot be solved exactly on the Real RAM.
We therefore need to relax the goal and ask for an approximation algorithm. Geometric Median has a very fast $(1+\varepsilon)$-approximation algorithm running in time $\mathcal{O}(nd\log^{3}(1/\varepsilon))$ [CohenLMPS16], so the reduction from Geometric Median to $\mathrm{EMDuT}_{2}$ does not rule out very fast approximation algorithms for $\mathrm{EMDuT}_{2}$.
@@ 3 Algorithm in One Dimension
For any $B,R\subset\mathbb{R}$ there is an optimal matching $\phi$ that is monotonically increasing.
@@
We say that $(b,b^{\prime})\in B\times B$ forms a crossing in a matching $\phi$ if $b>b^{\prime}$ and $\phi(b)<\phi(b^{\prime})$. Let $\phi$ be an optimal matching of $B,R\subseteq\mathbb{R}$ with the minimal number of crossings. If $\phi$ does not have any crossing, it is monotonically increasing. Hence, for the sake of contradiction assume that $(b,b^{\prime})$ is a crossing in $\phi$. Let $r=\phi(b^{\prime})$ and $r^{\prime}=\phi(b)$ and consider a matching $\phi^{\prime}$ that has $\phi^{\prime}(b)=r$ and $\phi^{\prime}(b^{\prime})=r^{\prime}$, and $\phi^{\prime}(x)=\phi(x)$ for every $x\in B\setminus\{b,b^{\prime}\}$.
We will show that $\mathcal{D}_{B,R}(\phi,0)\geq\mathcal{D}_{B,R}(\phi^{\prime},0)$. Combined with the fact that $\phi^{\prime}$ has less crossings than $\phi$, this yields a contradiction to the choice of $\phi$ as the optimal matching with the minimal number of crossings.
Note that $\mathcal{D}_{B,R}(\phi,0)\geq\mathcal{D}_{B,R}(\phi^{\prime},0)$ is equivalent to
$|b-r^{\prime}|+|b^{\prime}-r|\geq|b-r|+|b^{\prime}-r^{\prime}|.$ (3.1)
Since $b^{\prime}<b$ and $r^{\prime}<r$, inequality (3.1) follows from the fact below (by setting $x=b^{\prime}-r$, $\alpha=b-b^{\prime}$, and $\beta=r-r^{\prime}$). ∎
Fact 3.2.
For every $x\in\mathbb{R}$ and $\alpha,\beta>0$ it holds that $|x|+|x+\alpha+\beta|\geq|x+\alpha|+|x+\beta|$.
@@ 3.1 Symmetric Case
In the symmetric case ($|R|=|B|$), Lemma 3.1 uniquely defines an optimal matching. Let $B=\{b_{1},\ldots,b_{n}\}$ and $R=\{r_{1},\ldots,r_{n}\}$ be the points in increasing order. Now, the optimal translation $\tau^{*}$ is the value for $\tau$ that minimizes $\mathcal{D}_{B,R}(\phi,\tau)=\sum_{i=1}^{n}|b_{i}-r_{i}+\tau|$. Thus, it corresponds to the median of $b_{1}-r_{1},\ldots,b_{n}-r_{n}$, which we can compute in $\mathcal{O}(n\log n)$ time.
@@ 3.2 Asymmetric Case
In this section, we present an $\mathcal{O}(mn(\log n+\log^{2}m))$ time algorithm to compute $\mathrm{EMDuT}(B,R)$, for the case that $m\leq n$. Consider the cost $f(\tau)=\min_{\phi\in\Phi}\mathcal{D}_{B,R}(\phi,\tau)$ as a function of $\tau$. The minimum of this function is $\mathrm{EMDuT}(B,R)$. The main idea is then to sweep over the domain of $f$, increasing $\tau$ from -$\infty$ to $\infty$, while maintaining (a representation of) $f$ and a matching $\phi$ that realizes cost $f(\tau)=\mathcal{D}_{B,R}(\phi,\tau)$. We also maintain the best translation $\tau^{*}\leq\tau$ (i.e. with minimal cost) among the translations considered so far (and if there are multiple such translations, the smallest one), so at the end of our sweep, $\tau^{*}$ is thus an optimal translation.
@@
By Lemma 3.1, for any $\tau$, there exists an optimal monotonically increasing matching between $B+\tau$ and $R$. So, we restrict our attention to such monotonically increasing matchings. Observe that any such matching $\phi$ corresponds to a partition of $B$ into runs, i.e. maximal subsequences of consecutive points, $B_{1},\ldots,B_{z}$, so that the points $b_{t-k},\ldots,b_{t}$ in a run $B_{i}$ are matched to consecutive red points $r_{u-k},\ldots,r_{u}$, for some $r_{u}=\phi(b_{t})$. Moreover, for any such a matching $\phi$, the function $\mathcal{D}_{B,R}(\phi,\tau)$ is piecewise linear in $\tau$, and each breakpoint is a translation $\tau$ for which there is a pair $(b,r)\in B\times R$ with $b+\tau=r$. It then follows that $f(\tau)$ is also piecewise linear in $\tau$. Furthermore, the breakpoints of $f$ are of two types. A type (i) breakpoint is a translation such that there is a pair $(b,r)\in B\times R$ with $b+\tau=r$, and a type (ii) breakpoint if there are two different matchings $\phi,\phi^{\prime}$ that both realize the same minimum cost $\mathcal{D}_{B,R}(\phi,\tau)=\mathcal{D}_{B,R}(\phi^{\prime},\tau)$. We show the following key lemma, which lets us characterize the breakpoints of type (ii) more precisely.
Figure 3.1: Schematic representation of the graph $G=\phi\oplus\phi^{\prime}$ used in the proof of Lemma˜3.4. Each edge exists if and only if exactly one edge from either $\phi$ or $\phi^{\prime}$ is present. Green edges arise from the matching $\phi^{\prime}$, while yellow edges arise from the matching $\phi$.
@@
Let $\phi$ be an optimal monotone matching of $\mathrm{EMD}_{p}(B+\tau,R)$, and let $\phi^{\prime}$ be an optimal monotone matching of $\mathrm{EMD}_{p}(B+\tau^{\prime},R)$ for some $\tau^{\prime}>\tau$. Then, $\phi^{\prime}(b)\geq\phi(b)$ for all $b\in B$.
@@
We assume that all the points are distinct, as otherwise we can perturb them infinitesimally to resolve ties. Consider a bipartite graph $G$ defined as follows: the vertices of $G$ are $B\uplus R$, and we add an edge $(b,r)\in B\times R$ to $G$ if exactly one of the following conditions holds: (i) $\phi(b)=r$ and $\phi^{\prime}(b)\neq r$, or (ii) $\phi^{\prime}(b)=r$ and $\phi(b)\neq r$. The graph $G$ can be thought of as the exclusive-or of the matchings $\phi$ and $\phi^{\prime}$, see Figure˜3.1. We will now demonstrate that the connected components of this graph are paths. Then, considering that the matchings are monotone, it follows that the edges of these paths are non-crossing. This implies that consecutive red vertices on these paths are monotone. The lemma statement then easily follows.
@@
Observe that the maximum degree of graph $G$ is $2$, so the connected components of $G$ consist of cycles and paths. Assume that $C$ is a cycle. In that case, however, there exists a pair of edges from either $\phi$ or $\phi^{\prime}$ that intersect (see Figure˜3.2). This contradicts the assumption about the monotonicity of both $\phi$ and $\phi^{\prime}$. Hence, $C$ is not a cycle, and the proof of the claim follows. ∎
@@
Every vertex $b\in B$ has degree $2$ in $G$ unless $b$ is an isolated vertex. However, both vertices $p_{1}$ and $p_{\ell}$ are endpoints of a path, which means that $p_{1}$ and $p_{\ell}$ have degree $1$ in $G$. This means that $p_{1},p_{\ell}\notin B$. ∎
@@
In particular, there exists an index $k\in[\ell-2]$ such that $t_{k}\cdot t_{k+1}<0$. This means that either (a) $p_{k}>p_{k+2}$ and $p_{k+1}<p_{k+3}$, or (b) $p_{k}<p_{k+2}$ and $p_{k+1}>p_{k+3}$. In both of these cases, the intervals $(\min\{p_{k},p_{k+1}\},\max\{p_{k},p_{k+1}\})$ and $(\min\{p_{k+2},p_{k+3}\},\max\{p_{k+2},p_{k+3}\})$ intersect (see Figure˜3.3), which contradicts the assumption about the monotonicity of $\phi$ and $\phi^{\prime}$. ∎
Figure 3.3: Two cases of Claim˜3.7 in which a crossing occurs.
Now, we continue the proof of Lemma˜3.4. For the sake of contradiction, assume that $\phi^{\prime}(b)<\phi(b)$ for some $b\in B$. By Claim˜3.7, there exists a connected component $C$ of $G$ such that $\phi^{\prime}(b)<\phi(b)$ for every $b\in C\cap B$. Fix any such component $C$ and for every $b\in B$ let:
$\psi(b)\coloneq\begin{cases}\phi^{\prime}(b)&\text{ if }b\in C\cap B,\\ \phi(b)&\text{ otherwise.}\end{cases}$ and $\psi^{\prime}(b)\coloneq\begin{cases}\phi(b)&\text{ if }b\in C\cap B,\\ \phi^{\prime}(b)&\text{ otherwise.}\end{cases}$
Notice that $\psi$ and $\psi^{\prime}$ are both matchings of $\mathrm{EMD}_{p}(B+\tau,R)$ and $\mathrm{EMD}_{p}(B+\tau^{\prime},R)$ respectively. Now, recall that the cost of each matching is:
$\mathcal{D}_{B,R}(\phi,\tau)=$ $\sum_{b\in C\cap B}\lvert b+\tau-\phi(b)\rvert+\sum_{b\in B\setminus C}\lvert b+\tau-\phi(b)\rvert$,
$\mathcal{D}_{B,R}(\phi^{\prime},\tau^{\prime})=$ $\sum_{b\in C\cap B}\lvert b+\tau^{\prime}-\phi^{\prime}(b)\rvert+\sum_{b\in B\setminus C}\lvert b+\tau^{\prime}-\phi^{\prime}(b)\rvert$,
$\mathcal{D}_{B,R}(\psi,\tau)=$ $\sum_{b\in C\cap B}\lvert b+\tau-\phi^{\prime}(b)\rvert+\sum_{b\in B\setminus C}\lvert b+\tau-\phi(b)\rvert,\text{ and }$
$\mathcal{D}_{B,R}(\psi^{\prime},\tau^{\prime})=$ $\sum_{b\in C\cap B}\lvert b+\tau^{\prime}-\phi(b)\rvert+\sum_{b\in B\setminus C}\lvert b+\tau^{\prime}-\phi^{\prime}(b)\rvert$.